Two parallel current carrying conductor exert some force on each other.If the current in both the conductors is doubled,the force becomes k times.find the value of k
1
Expert's answer
2012-11-01T08:51:14-0400
According to the Ampere's force law the force of interaction between two parallel current carrying conductors is

F1 = 2μI²/(4πr)

before doubling and

F2 = 2μ(2I)²/(4πr)

after doubling. Therefore, the coefficient k is

k = F2/F1 = 2μ(2I)²/(4πr) / (2μI²/(4πr)) = (2I)²/(I²) = 4.